### v1

@@ 4.1 AMPDA and its interpretations
where ${(I)}$ follows from the definition of $\widetilde{F}$ (see (13)), ${(II)}$ follows from $f\geq 0$ and $y^{k}\in\partial g(x^{k})$, ${(III)}$ follows from the Lipschitz continuity of $f$ and the Cauchy-Schwarz Inequality, and ${(IV)}$ holds since the Lipschitz continuity of $\nabla h_{1}$ on $\mathcal{B}(x^{k},\delta_{x})$ and $z^{k}\in\partial h_{2}(x^{k})$ imply that
$h(x)=h_{1}(x)-h_{2}(x)$
$\leq h_{1}(x^{k})-h_{2}(x^{k})+\left\langle\nabla h_{1}(x^{k})-z^{k},x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}$
$=h(x^{k})+\left\langle\nabla h_{1}(x^{k})-z^{k},x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}.$
Next, in the minimization stage one naturally updates $x$ by solving $\min\{\mathcal{M}(x\mid x^{k},c_{k}):x\in\mathbb{R}^{n}\}$. However, as the Lipschitz moduli $L_{f,k}$ and $L_{\nabla h_{1},k}$ are generally difficult to estimate, we use $1/(2\alpha)$ with $\alpha>0$ to replace the unknown term $c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}$ in $\mathcal{M}(x\mid x^{k},c_{k})$, where $\alpha$ is determined by a backtracking line search procedure. Invoking this, (35) and the fact that $c_{k}=1/g(x^{k})$, we actually in the minimization stage construct $x^{k+1}$ via
(36) $x^{k+1}\in\arg\min\Big{\{}c_{k}f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\frac{1}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\Big{\}}.$
@@
Third, the line search procedure is stopped once $\widehat{x}^{k}\in\Omega$ and (34) is satisfied. Note that $(y^{k},z^{k})\in\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})$ holds for all $k\in\mathbb{N}$, owing to $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$. Invoking this and (32), we immediately see that the condition (34) is generally stronger than the sufficient descent condition of $F$ at $\widehat{x}^{k}$, that is,
(37) $F(\widehat{x}^{k})+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k}).$
Although the subsequential convergence can be still guaranteed if the terminating condition (34) of the line search scheme in AMPDA is replaced by (37), it is generally difficult to show the convergence of the full solution sequence generated by this altered algorithm. In contrast, we can not only establish the subsequential convergence but also the convergence of the entire solution sequence generated by the AMPDA equipped with (34) under the KL assumption.
@@ 4.2 Convergence analysis for the AMPDA
In this subsection, we conduct convergence analysis for the AMPDA. To this end, we first introduce the following basic assumption concerning its initial point.
@@
Note that the boundedness of the level set associated with the extended objective is a very standard assumption in nonconvex optimization, while its closeness automatically holds if the extended objective is closed. Nevertheless, the $F$ defined in (12) may be not closed, which necessitates the closedness condition in 2. In view of [44, Proposition 4.4] and the closedness of $h$ and $\mathcal{C}$, we see that $F$ is closed on $\mathbb{R}^{n}$ if $f$ and $g$ do not attain zero simultaneously. In the case where $f$ and $g$ can attain zero simultaneously, we can still guarantee the closedness of $\mathcal{X}$ via appropriately selecting $x^{0}$, as shown in the following proposition.
Proposition 4.1.
@@
(38) $F(x^{0})<\inf\left\{\mathop{\lim\inf}_{z\to x}F(z):x\in\mathcal{O}\right\}.$
Proof 4.2.
@@
(39) $f(x^{\star})+h(x^{k})g(x^{\star})$ $\leq F(x^{0})g(x^{k})+h(x^{k})(g(x^{\star})-g(x^{k}))$
$\leq F(x^{0})g(x^{k})+\sup\{|h(x^{k})|:k\in\mathbb{N}\}|g(x^{\star})-g(x^{k})|.$
@@
(40) $f(x^{\star})+h(x^{\star})g(x^{\star})\leq F(x^{0})g(x^{\star}),$
owing to the closed $f$, the closed $g(x^{\star})h$, and the continuous $g$. From (40) and the closedness of $\mathcal{C}$, one can derive $x^{\star}\in\mathcal{X}$ if $g(x^{\star})>0$. In fact, $g(x^{\star})>0$ is guaranteed when $x^{0}\in\mathrm{dom}(F)$ satisfies (38). Assume on the contrary that (38) holds but $g(x^{\star})=0$, which forces $f(x^{\star})=0$ according to (40) and $f\geq 0$. In other words, $g(x^{\star})=0$ leads to $x^{\star}\in\mathcal{O}$. This, along with (38), implies that $F(x^{0})<\mathop{\lim\inf}_{k\to\infty}F(x^{k})$, which contradicts the fact that $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$.
We remark that for both problems (2) and (3), the zero vector is the unique point at which both $f$ and $g$ vanish. In view of Proposition 4.1 and these two problems, we deduce that if $x^{0}\in\mathbb{R}^{n}$ satisfies $F(x^{0})<\mathop{\lim\inf}_{z\to 0}F(z)=1+\frac{\lambda}{2}\|b-\mathcal{T}_{\mu}(b)\|_{2}^{2}$, then the level set $\mathcal{X}$ is closed. This together with the boundedness of the constraint $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$ shows that 2 is fulfilled for problems (2) and (3) if the initial point $x^{0}$ is properly chosen such that
(41) $F(x^{0})<1+\frac{\lambda}{2}\|b-\mathcal{T}_{\mu}(b)\|_{2}^{2}.$
Now with the help of 2, we can establish the following technical lemma, which will be frequently used in our subsequent convergence analysis.
Lemma 4.3.
Suppose that 2 holds. Then, we have $m_{g}:=\inf\{g(x):x\in\mathcal{X}\}>0$. Moreover, there exists $\Delta>0$ such that the following statements hold with the set $\mathcal{X}_{\Delta}$ defined by
@@
Proof 4.4.
Since the continuous $g$ is strictly positive on the compact $\mathcal{X}$, we deduce that $m_{g}>0$ and the statement (i) holds. Then, $f$, $g$, and $\nabla h_{1}$ are globally Lipschitz continuous on $\mathcal{X}_{\Delta}$, since they are locally Lipschitz continuous around each point in the compact $\mathcal{X}_{\Delta}\subseteq\Omega$ ([13, Chapter 1, 7.5 Exercise (c)]). This, along with the statement (i), implies that the statement (ii) holds. Lastly, $M_{\partial g}$, $M_{f/g^{2}}$, and $M_{g}$ are finite due to the boundedness of $\cup_{x\in\mathcal{X}}\partial g(x)$, continuity of $f/g^{2}$ on the compact $\mathcal{X}_{\Delta}$, and the continuity of $g$ on the compact $\mathcal{X}$, respectively. This completes the proof.
@@
Proposition 4.5.
Suppose that 2 holds. Then the AMPDA is well-defined. Specifically, there exists $\underline{\alpha}_{\sigma}>0$ such that, for all $k\in\mathbb{N}$, Step 2 of AMPDA terminates at some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. Moreover, the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}\subseteq\Omega\cap\mathcal{C}$ and satisfies
(42) $F(x^{k+1})+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}_{2}\leq F(x^{k}).$
Proof 4.6.
@@
(43) $\alpha c_{k}f(\widehat{x}^{k})+\frac{1}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}+\left\langle\widehat{x}^{k}-x^{k},\alpha(\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k})\right\rangle\\ \leq\alpha c_{k}f(x^{k}).$
@@
(44) $\|\widehat{x}^{k}-x^{k}\|_{2}\leq\alpha|g(x^{k})|\left\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\right\|_{2}\\ +\sqrt{\alpha^{2}|g(x^{k})|^{2}\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\|_{2}^{2}+2\alpha f(x^{k})}.$
@@
(45) $\|\widehat{x}^{k}-x^{k}\|_{2}\leq\Delta/2,\text{~{}~{}~{}~{}~{}for any $\alpha\in(0,\underline{\alpha}_{\Delta}]$}.$
@@
Next, we shall show that there exists $\underline{\alpha}_{\sigma}\in(0,\underline{\alpha}_{\Delta}]$ such that $\widehat{x}^{k}\in\Omega$ and satisfies (34) stated in Step 2 of AMPDA whenever $\alpha\in(0,\underline{\alpha}_{\sigma}]$. Let $\alpha\in(0,\underline{\alpha}_{\Delta}]$. Clearly, with the help of Lemma 4.3 (i), the statement that $\widehat{x}^{k}\in\Omega$ holds due to $\widehat{x}^{k}\in\mathcal{X}_{\Delta}$. Besides, from (45), we have $\widehat{x}^{k}\in\mathcal{B}(x^{k},\Delta)$. This, together with the Lipschitz continuity of $\nabla h_{1}$ on $\mathcal{B}(x^{k},\Delta)\subseteq\mathcal{X}_{\Delta}$ (see Lemma 4.3 (ii)), implies that
(46) $h_{1}(\widehat{x}^{k})\leq h_{1}(x^{k})+\left\langle{\nabla h_{1}(x^{k})},{\widehat{x}^{k}-x^{k}}\right\rangle+\frac{L_{\nabla h_{1}}}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2},$
@@
(47) $c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},z^{k}+c^{2}_{k}f(x^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}\right)\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq c_{k}f(x^{k})+h_{1}(x^{k}).$
@@
(48) $c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},c^{2}_{k}f(x^{k})y^{k}\right\rangle$
@@
where ${(I)}$ follows from the Cauchy-Schwarz Inequality and Lemma 4.3 (ii) with the positive $L_{g}$ and $L_{f/g^{2}}$ being the Lipschitz moduli of $g$ and $f/g^{2}$ on $\mathcal{X}_{\Delta}$, respectively, and ${(II)}$ follows from Lemma 4.3 (iii). Combining (47) and (48), we derive that
(49) $\hat{c}^{2}f(\widehat{x}^{k})\left(2g\left(\widehat{x}^{k}\right)-g(x^{k})\right)+\left\langle x^{k}-\widehat{x}^{k},z^{k}+\hat{c}^{2}f(\widehat{x}^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}-M_{f/g^{2}}L_{g}^{2}-L_{f/g^{2}}M_{\partial g}M_{g}\right)\left\|\widehat{x}^{k}-x^{k}\right\|_{2}^{2}\leq c_{k}f(x^{k})+h_{1}(x^{k}).$
@@
(50) $Q(x^{k+1},y^{k},z^{k},1/g(x^{k+1}))+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|_{2}^{2}\leq F(x^{k}),$
which, together with (32), indicates (42). In view of (42) and $x^{k}\in\mathcal{X}$, we obtain $x^{k+1}\in\mathcal{X}$. Consequently, we conclude this proposition by induction.
Proposition 4.5 and the inequality (42) immediately lead to the next corollary.
Corollary 4.7.
Suppose that 2 holds. Then, the $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ generated by AMPDA is bounded. Besides, the limit $F_{\infty}:=\lim_{k\to\infty}F(x^{k})$ exists, and it holds that
(51) $\lim_{k\to\infty}\|x^{k+1}-x^{k}\|_{2}=0.$
Proof 4.8.
Firstly, since $x^{k}\in\mathcal{X}$ and the functions $g$ and $h_{2}$ are real-valued convex, we derive $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ is bounded due to $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$ for all $k\in\mathbb{N}$, $c_{k}=1/g(x^{k})$ and the compact $\mathcal{X}\subseteq\Omega$. Secondly, since $F$ is continuous on $\Omega\cap\mathcal{C}\supseteq\mathcal{X}$, the function $F$ is bounded below on the compact $\mathcal{X}$, which together with (42) indicates that the limit $F_{\infty}:=\lim_{k\to\infty}F(x^{k})$ exists. Thirdly, the equation (51) is derived by passing to the limit on the both sides of (42) with $k\to\infty$.

### v2

@@ 4.1 AMPDA and its interpretations
where (I){(I)} follows from the definition of $\widetilde{F}$ (see (10)); (II){(II)} follows from $f\geq 0$ and $y^{k}\in\partial g(x^{k})$; (III){(III)} follows from the Lipschitz continuity of $f$ and the Cauchy-Schwarz Inequality, and (IV){(IV)} follows from the inequality $h(x)\leq h(x^{k})+\left\langle\nabla h_{1}(x^{k})-z^{k},x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}$. This inequality holds because the convexity of $h_{2}$ implies $h_{2}(x)\geq h_{2}(x^{k})+\left\langle z^{k},x-x^{k}\right\rangle$, while the Descent Lemma applied to the locally Lipschitz differentiable function $h_{1}$ yields $h_{1}(x)\leq h_{1}(x^{k})+\left\langle\nabla h_{1}(x^{k}),x-x^{k}\right\rangle+\frac{L_{\nabla h_{1},k}}{2}\|x-x^{k}\|_{2}^{2}$. Next, in the minimization stage one naturally updates $x$ by solving $\min\{\mathcal{M}(x\mid x^{k},c_{k}):x\in\mathbb{R}^{n}\}$. However, as the Lipschitz moduli $L_{f,k}$ and $L_{\nabla h_{1},k}$ are generally difficult to estimate, we use $1/(2\alpha)$ with $\alpha>0$ to replace the unknown term $c_{k}^{2}L_{f,k}\|y^{k}\|_{2}+\frac{L_{\nabla h_{1},k}}{2}$ in $\mathcal{M}(x\mid x^{k},c_{k})$, where $\alpha$ is determined by a backtracking line search procedure. Invoking this, (36) and the fact that $c_{k}=1/g(x^{k})$, we actually in the minimization stage construct $x^{k+1}$ via
$x^{k+1}\in\arg\min\Big\{c_{k}f(x)+\left\langle\nabla h_{1}(x^{k})-c_{k}^{2}f(x^{k})y^{k}-z^{k},x-x^{k}\right\rangle\\ +\frac{1}{2\alpha}\|x-x^{k}\|_{2}^{2}:x\in\mathcal{C}\Big\}.$ (37)
@@
Third, the line search procedure is stopped once $\widehat{x}^{k}\in\Omega$ and (35) is satisfied. Using the fact that $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$ and the Fenchel-Young equality, it is easy to verify that
$Q(\widehat{x}^{k},y^{k},z^{k},1/g(\widehat{x}^{k}))=\frac{f(\widehat{x}^{k})}{g^{2}(\widehat{x}^{k})}(2g(\widehat{x}^{k})-g(x^{k}))+h_{1}(\widehat{x}^{k})-h_{2}(x^{k})\\ +\left\langle{x^{k}-\widehat{x}^{k}},{z^{k}+\frac{f(\widehat{x}^{k})}{g^{2}(\widehat{x}^{k})}y^{k}}\right\rangle.$ (38)
Hence, in practice we do not need to evaluate the conjugate terms $g^{*}$ and $h^{*}_{2}$ in the line-search procedure. Also, invoking $(y^{k},z^{k})\in\mathrm{dom}(g^{*})\times\mathrm{dom}(h_{2}^{*})$ and (33), we immediately see that the condition (35) is generally stronger than the sufficient descent condition of $F$ at $\widehat{x}^{k}$, that is,
$F(\widehat{x}^{k})+\frac{\sigma}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq F(x^{k}).$ (39)
Although the subsequential convergence can be still guaranteed if the terminating condition (35) of the line search scheme in AMPDA is replaced by (39), it is difficult to show the convergence of the full solution sequence generated by this altered algorithm under the KL assumption on $F$. This is because it is intractable to verify that $F$ satisfies the relative error condition (condition (ii) in Proposition 2.1) at the solution sequence without additional smooth assumption on $g$. In contrast, we can use $Q$ defined in (32) as an auxiliary function, and prove that it generally satisfies conditions (i)-(iii) of Proposition 2.1 at the sequence $\{(x^{k+1},y^{k},z^{k},c_{k+1}):k\in\mathbb{N}\}$ generated by AMPDA equipped with the line-search condition (35), as shown in Section 5. Due to these and invoking Proposition 2.1, we immediately obtain the sequential convergence of AMPDA under the KL assumption on $Q$.
@@ 4.2 Convergence analysis for the AMPDA
In this subsection, we conduct convergence analysis for AMPDA. To this end, we first introduce the following basic assumption concerning its initial point.
@@
Assumption 2 requires that the set $\mathcal{X}$ is bounded and closed. The boundedness of the level set associated with the extended objective is a very standard assumption in nonconvex optimization. In particular, for problem (1), the boundedness of $\mathcal{X}$ is automatically guaranteed if the constraint set $\mathcal{C}$ is bounded or the function $h$ is coercive (i.e., $\lim_{\|x\|_{2}\to\infty}h(x)=+\infty$). However, the closedness of $\mathcal{X}$ often requires a more careful discussion since $F$ is probably not closed. In view of (NaZhang-QiaLi:2022SIAM-OPT, , Proposition 4.4), and using the closedness of $h$ and $\mathcal{C}$, the objective $F$ is a closed function whenever $f$ and $g$ do not attain zero simultaneously. Consequently, any level set of $F$, including $\mathcal{X}$, is closed. In the case where $f$ and $g$ can attain zero simultaneously, we can still guarantee the closedness of $\mathcal{X}$ via appropriately selecting $x^{0}$, as shown in the following proposition.
Proposition 4.1
@@
$F(x^{0})<\inf\left\{\mathop{\lim\inf}_{z\to x}F(z):x\in\mathcal{O}\right\}.$ (40)
Proof
@@
$f(x^{k})+h(x^{k})g(x^{\star})\leq F(x^{0})g(x^{k})+h(x^{k})(g(x^{\star})-g(x^{k}))\\ \leq F(x^{0})g(x^{k})+\sup\{|h(x^{k})|:k\in\mathbb{N}\}|g(x^{\star})-g(x^{k})|.$ (41)
@@
$f(x^{\star})+h(x^{\star})g(x^{\star})\leq F(x^{0})g(x^{\star}),$ (42)
owing to the closed $f$, the closed $g(x^{\star})h$, and the continuous $g$. From (42) and the closedness of $\mathcal{C}$, one can derive $x^{\star}\in\mathcal{X}$ if $g(x^{\star})>0$. In fact, $g(x^{\star})>0$ is guaranteed when $x^{0}\in\mathrm{dom}(F)$ satisfies (40). Assume on the contrary that (40) holds but $g(x^{\star})=0$, which forces $f(x^{\star})=0$ according to (42) and $f\geq 0$. In other words, $g(x^{\star})=0$ leads to $x^{\star}\in\mathcal{O}$. This, along with (40), implies that $F(x^{0})<\mathop{\lim\inf}_{k\to\infty}F(x^{k})$, which contradicts $\{x^{k}:k\in\mathbb{N}\}\subseteq\mathcal{X}$. $\square$
We remark that for both problems (2) and (3), the zero vector is the unique point at which both $f$ and $g$ vanish. In view of Proposition 4.1 and these two problems, we deduce that if $x^{0}\in\mathbb{R}^{n}$ satisfies
$F(x^{0})<\mathop{\lim\inf}_{z\to 0}F(z)=1+\frac{\lambda}{2}\|b-\mathcal{T}_{\mu}(b)\|_{2}^{2},$ (43)
then the level set $\mathcal{X}$ is closed. This together with the boundedness of the constraint set $\{x\in\mathbb{R}^{n}:\underline{x}\leq x\leq\overline{x}\}$ shows that Assumption 2 is fulfilled for problems (2) and (3) if the initial point $x^{0}$ is properly chosen. Moreover, as we shall demonstrate in Section 6, such an initial point $x^{0}$ satisfying (43) can be easily constructed in practice. At this stage, however, we concentrate on the convergence analysis of the proposed algorithm.
With the aid of Assumption 2, we now establish the following technical lemma, which will be frequently used in our subsequent analysis.
Lemma 4.2
Suppose that Assumption 2 holds. Then, we have $m_{g}:=\inf\{g(x):x\in\mathcal{X}\}>0$. Moreover, there exists $\Delta>0$ such that the following statements hold with the set $\mathcal{X}_{\Delta}$ defined by
@@
Proof
Since the continuous $g$ is strictly positive on the compact $\mathcal{X}$, we deduce that $m_{g}>0$ and the statement (i) holds. Then, $f$, $g$, and $\nabla h_{1}$ are globally Lipschitz continuous on $\mathcal{X}_{\Delta}$, since they are locally Lipschitz continuous around each point in the compact $\mathcal{X}_{\Delta}\subseteq\Omega$ ((Clarke:2008BookNonsmooth_analysis_and_control_theory, , Chapter 1, 7.5 Exercise (c))). This, along with the statement (i), implies that the statement (ii) holds. Lastly, $M_{\partial g}$, $M_{f/g^{2}}$, and $M_{g}$ are finite due to the boundedness of $\cup_{x\in\mathcal{X}}\partial g(x)$ ((Bertsekas:Convexoptimizationtheory, , Proposition 5.4.2)), continuity of $f/g^{2}$ on the compact $\mathcal{X}_{\Delta}$, and the continuity of $g$ on the compact $\mathcal{X}$, respectively. This completes the proof. $\square$
@@
Proposition 4.3
Suppose Assumption 2 holds. Then AMPDA is well-defined. Specifically, there exists $\underline{\alpha}_{\sigma}>0$ such that, for all $k\in\mathbb{N}$, Step 2 of AMPDA terminates at some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. Moreover, the sequence $\{x^{k}:k\in\mathbb{N}\}$ generated by AMPDA falls into the level set $\mathcal{X}\subseteq\Omega\cap\mathcal{C}$ and satisfies
$F(x^{k+1})+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|^{2}_{2}\leq F(x^{k}),$ (44)
where $\sigma>0$ is given in Step 0 of AMPDA.
Proof
@@
$\alpha c_{k}f(\widehat{x}^{k})+\frac{1}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2}+\left\langle\widehat{x}^{k}-x^{k},\alpha(\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k})\right\rangle\\ \leq\alpha c_{k}f(x^{k}).$ (45)
@@
$\|\widehat{x}^{k}-x^{k}\|_{2}\leq\alpha|g(x^{k})|\left\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\right\|_{2}\\ +\sqrt{\alpha^{2}|g(x^{k})|^{2}\|\nabla h_{1}(x^{k})-z^{k}-c^{2}_{k}f(x^{k})y^{k}\|_{2}^{2}+2\alpha f(x^{k})}.$ (46)
@@
$\|\widehat{x}^{k}-x^{k}\|_{2}\leq\Delta/2,\text{~~~~~for any $\alpha\in(0,\underline{\alpha}_{\Delta}]$}.$ (47)
@@
Next, we shall show that there exists $\underline{\alpha}_{\sigma}\in(0,\underline{\alpha}_{\Delta}]$ such that $\widehat{x}^{k}\in\Omega$ and satisfies (35) stated in Step 2 of AMPDA whenever $\alpha\in(0,\underline{\alpha}_{\sigma}]$. Let $\alpha\in(0,\underline{\alpha}_{\Delta}]$. Clearly, with the help of Lemma 4.2 (i), the statement that $\widehat{x}^{k}\in\Omega$ holds due to $\widehat{x}^{k}\in\mathcal{X}_{\Delta}$. From (47), we have $\widehat{x}^{k}\in\mathcal{B}(x^{k},\Delta)$. This, together with the Lipschitz continuity of $\nabla h_{1}$ on $\mathcal{B}(x^{k},\Delta)\subseteq\mathcal{X}_{\Delta}$ (see Lemma 4.2 (ii)), implies that
$h_{1}(\widehat{x}^{k})\leq h_{1}(x^{k})+\left\langle{\nabla h_{1}(x^{k})},{\widehat{x}^{k}-x^{k}}\right\rangle+\frac{L_{\nabla h_{1}}}{2}\|\widehat{x}^{k}-x^{k}\|_{2}^{2},$ (48)
@@
$c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},z^{k}+c^{2}_{k}f(x^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}\right)\|\widehat{x}^{k}-x^{k}\|_{2}^{2}\leq c_{k}f(x^{k})+h_{1}(x^{k}).$ (49)
@@
$c_{k}f(\widehat{x}^{k})+\left\langle x^{k}-\widehat{x}^{k},c^{2}_{k}f(x^{k})y^{k}\right\rangle$ (50)
@@
where (I){(I)} follows from the Cauchy-Schwarz Inequality and Lemma 4.2 (ii) with the positive $L_{g}$ and $L_{f/g^{2}}$ being the Lipschitz moduli of $g$ and $f/g^{2}$ on $\mathcal{X}_{\Delta}$, respectively, and (II){(II)} follows from Lemma 4.2 (iii). Combining (49) and (50), we derive that
$\hat{c}^{2}f(\widehat{x}^{k})\left(2g\left(\widehat{x}^{k}\right)-g(x^{k})\right)+\left\langle x^{k}-\widehat{x}^{k},z^{k}+\hat{c}^{2}f(\widehat{x}^{k})y^{k}\right\rangle+h_{1}(\widehat{x}^{k})\\ +c_{k}\left(\frac{1}{2\alpha}-\frac{L_{\nabla h_{1}}}{2}-M_{f/g^{2}}L_{g}^{2}-L_{f/g^{2}}M_{\partial g}M_{g}\right)\left\|\widehat{x}^{k}-x^{k}\right\|_{2}^{2}\\ \leq c_{k}f(x^{k})+h_{1}(x^{k}).$ (51)
@@
$Q(x^{k+1},y^{k},z^{k},1/g(x^{k+1}))+\frac{\sigma}{2}\|x^{k+1}-x^{k}\|_{2}^{2}\leq F(x^{k}),$ (52)
which, together with (33), indicates (44). In view of (44) and $x^{k}\in\mathcal{X}$, we obtain $x^{k+1}\in\mathcal{X}$. Consequently, we conclude this proposition by induction. $\square$
Proposition 4.3 and the inequality (44) immediately lead to the next corollary. In particular, we prove that the AMPDA can find an $\epsilon$-critical point of $F$ within $\mathcal{O}(\epsilon^{-2})$ iterations.
Corollary 4.4
Suppose that Assumption 2 holds. Then the following statements hold.
  1. (i)
    The sequence $\{(x^{k},y^{k},z^{k},c_{k}):k\in\mathbb{N}\}$ generated by AMPDA is bounded.
  2. (ii)
    The sequence $\{F(x^{k}):k\in\mathbb{N}\}$ descends monotonically with $F_{\infty}:=\inf\{F(x^{k}):k\in\mathbb{N}\}=\lim_{k\to\infty}F(x^{k})$, and we have
    $\lim_{k\to\infty}\|x^{k+1}-x^{k}\|_{2}=0.$ (53)
  3. (iii)
    Given $\epsilon>0$, there exists a $\hat{k}\leq K$ with
    $K:=\frac{2(F(x^{0})-F_{\infty})}{\sigma\epsilon^{2}},$
    such that $x^{\hat{k}}$ is an $\epsilon$-critical point of $F$.
Proof
First, we prove the statement (i). The boundedness of $\{(x^{k},c_{k}):k\in\mathbb{N}\}$ follows from the compactness of $\mathcal{X}$ and $c_{k}=1/g(x^{k})$. Also from the compactness of $\mathcal{X}$, we derive that $\{(y^{k},z^{k}):k\in\mathbb{N}\}$ is bounded, since $g$ and $h_{2}$ are real-valued convex and $(y^{k},z^{k})\in\partial g(x^{k})\times\partial h_{2}(x^{k})$. Next we prove the statement (ii). Since $F$ is continuous on $\Omega\cap\mathcal{C}\supseteq\mathcal{X}$, the function $F$ is bounded below on the compact $\mathcal{X}$. This along with (44) yields $F_{\infty}=\lim_{k\to\infty}F(x^{k})$. Then, by passing to the limit on the both sides of (44) with $k\to\infty$, we derive (53), and the statement (ii) follows. Finally, we prove the statement (iii). For all $K>0$, we have
$F(x^{K})+\frac{\sigma}{2}\sum_{k=0}^{K-1}\|x^{k+1}-x^{k}\|^{2}_{2}\leq F(x^{0}),$ (54)
by summing up (44) over k=0,1,,K1k=0,1,...,K-1. Invoking Proposition 4.3 and (34), we have
$x^{k+1}\in\mathrm{prox}_{\alpha_{k}c_{k}f_{\mathcal{C}}}\left(x^{k}-\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})\right),$ (55)
for some $\alpha_{k}\geq\min\{\underline{\alpha},\underline{\alpha}_{\sigma}\}\gamma$. From (54), (55) and $F_{\infty}\leq F(x^{K})$, we derive
$\min_{0\leq k\leq K-1}\mathrm{dist}(x^{k},\mathrm{prox}_{\alpha_{k}c_{k}f_{\mathcal{C}}}(x^{k}-\alpha_{k}(\nabla h_{1}(x^{k})-z^{k}-c_{k}^{2}f(x^{k})y^{k})))\\ \leq\sqrt{\frac{2}{\sigma K}(F(x^{0})-F_{\infty})}.$ (56)
According to Definition 5 and (56), we conclude the statement (iii). $\square$